OpenStudy (anonymous):

which is equal to 8^-10/8^5 ? A.8^50 B.8^15 C.8^5 D.8^2

OpenStudy (danjs):

\[\frac{ 8^{-10} }{ 8^5 }\]

OpenStudy (danjs):

recall \[\frac{ a^{-b} }{ 1 } = \frac{ 1 }{ a^b }\]

OpenStudy (danjs):

\[\frac{ 8^{-10} }{ 8^5 } = \frac{ 1 }{ 8^5*8^{10} }\]

OpenStudy (danjs):

none of the answers are right

OpenStudy (danjs):

\[8^a*8^b = 8^{a+b}\]

Nnesha (nnesha):

hmm there is a right answer :)